Answer this question for each part, (please provide detailed answers with Formula)
Problem 1: Jane Botosan operates a bed and breakfast hotel in a resort area near Lake Michigan. Depreciation on the hotel is S60,000 per year. Jane employs a maintenance person at an annual salary of $41,000 and a cleaning person at an annual salary of S24,000. Real estate taxes are $10,000 per year. The rooms rent at an average price of S60 per person per night including breakfast. Other costs are laundry and cleaning service at a cost of S10 per person per night and the cost of food which is S5 per person per night. Instructions (a) Determine the variable costs per person per night and the total annually fixed costs. (b) Determine the break-even point in (a) number of rentals and (b) in dollars using the contribution margin technique. (c) If the current number of rentals is 4,000, by what percentage can rentals decrease before Jane has to worry about having a net loss? That is, what is its margin of safety ratio? (d) Jane is considering upgrading the breakfast service to attract more business and increase prices. This will cost an additional $3 for food costs per person per night. Jane feels she can increase the room rate to $68 per person per night. Determine the number of rentals and the sales revenue Jane needs to break even if the changes are made.
Answer #1

A) VARIABLE COST

CLEANING SERVICE 10

COST OF FOOD 5

VARIABLE COST    =15

FIXED COST=DEPRECIATION+MAINTENANCE PERSON SALARY+CLEANING PERSON SALARY+REAL ESTATE TAXES

=60000+41000+24000+10000

=135000

B) BREAK EVEN (RENTALS)=135000/(0-15)=135000/45=3000

BREAK EVEN (DOLLARS)=135000*60/45=180000

C)CURRENT RENTALS=4000

BREAK EVEN RENTALS=3000

MARGIN OF SAFETY=4000-3000=1000

THEREFOE BY 25%(1000/4000*100) CAN RENTALS DECREASE BEFORE JANE HAS TO WORRY

D)BREAK EVEN (RENTALS)=135000/(68-18)=135000/50=2700

BREAK EVEN (SALES REVENUE)=135000*68/50=183600